The Royal Fruit Company produces two types of fruit drinks. The first type is 80% pure fruit juice, and the second type is 100% pure fruit juice. The company is
attempting to produce a fruit drink that contains 95% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 220
pints of a mixture that is 95% pure fruit juice?
First drink ? Pints
Second drink ? Pints

Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

80% pure fruit juice, be x pints
100% pure fruit juice required will be 220-x pints
Total Quantity =220 pints
100%x +80%(220-x) = 220 * 95%
100x +80(220-x) = 220 * 95
100x+17600-80x = 20900
20x = 3300
x = 165
80 % juice 165 pints
100% juice 55 pints

Answer by greenestamps(13200)



Here is a quick informal solution, if formal algebra is not required.

(1) 95% is 3/4 of the way from 80% to 100% (picture the three percentages 80, 95, and 100 on a number line, if it helps)
(2) Therefore 3/4 of the total 220 pints should be the 100% fruit juice

ANSWER: 3/4 of 220 pints, or 165 pints, of the 100% fruit juice; the other 55 pints of the 80% fruit juice.

CHECK: 1.00(165)+.80(55)=165+44=209
.95(220)=209